YOUR COMPANY CURRENTLY PRODUCES 10000 STOCK TANK BARRELS OF OIL PER DAY STBOPD THE QUOTED R/P RATIO FOR YOUR COMPANY IS 5 YEARS WHAT IS THE VOLUME OF RESERVES THAT YOUR COMPANY HAS CURRENTLY ON THE BOOKS HOW DOES IT COMPARE TO THE SAME RATIO FOR THE UNITED STATE AS A WHOLE IF YOU HAVE THE SAME AMOUNT OF ESTIMATED RESERVES BUT YOU PRODUCTION DROPS BY 10% PER YEAR CONSTRUCT A TABLE SHOWING THE REMAINING RESERVES VOLUMES FOR THE NEXT 20 YEARS ASSUME START DATE 12/31/2018
Answer #1

Solution :

1. Volume of reserves that company has currently in books :

R/P Ratio is known as Reserves to production per year Ratio which is used to calculate the remaining non renewable resources expressed in time. (Amount of known Reserve / Production per year)

Here we have given

R/P Ratio = 5 Times

Production per day = 10000 stock tanks barrels

Production per year = 10000*365 days = 3.65 million stock tanks barrels

Therefore ,

R/P Ratio =Amount of known reserve / Production per year

5 = Amount of known reserve/3.65

Amount of known reserve is 18.2 million stock tanks.

2. While comparing with the Country USA as a whole we need to compare the R/P Ratio of the company with the R/P Ratio of USA as a whole .

Fact from wikipedia : Proven oil reserves in the United States were 36.4 billion barrels (5.79×109 m3) of crude oil as of the end of 2014, excluding the Strategic Petroleum Reserve.

Therefore while comparing 18.2 Million stock tanks barrels with 36.4 Billion Barrels ,it comes out to be 0.05 % .

3. Table showing amount of reserves for next 20 Years

Year End

Reserves in Beginning (Stock tank Barrel /STB)

(A)

Production per year (decreasing 10% per annum)

(B)

Consumption per year

(C)

Reserve in year End

(A+B-C)

1 18500000 (18.5 Million STB) 3285000.00 n/a 21785000.00
2 21785000.00 2956500.00 n/a 24741500.00
3 24741500.00 2660850.00 n/a 27402350.00
4 27402350.00 2394765.00 n/a 29797115.00
5 29797115.00 2155288.50 n/a 31952403.50
6 31952403.50 1939759.65 n/a 33892163.15
7 33892163.15 1745783.69 n/a 35637946.84
8 35637946.84 1571205.32 n/a 37209152.15
9 37209152.15 1414084.78 n/a 38623236.94
10 38623236.94 1272676.31 n/a 39895913.24
11 39895913.24 1145408.68 n/a 41041321.92
12 41041321.92 1030867.81 n/a 42072189.73
13 42072189.73 927781.03 n/a 42999970.75
14 42999970.75 835002.92 n/a 43834973.68
15 43834973.68 751502.63 n/a 44586476.31
16 44586476.31 676352.37 n/a 45262828.68
17 45262828.68 608717.13 n/a 45871545.81
18 45871545.81 547845.42 n/a 46419391.23
19 46419391.23 493060.88 n/a 46912452.11
20 46912452.11 443754.79 n/a 47356206.90

Note :

A. We are assuming as there is no consumption from these reserves as if we do so the reserve balance will go negative in year 13th itself and it is not possible for reserves to go below 0.

B. Alternatively we can decrease consumption per year at the rate of 10% which will adjust the decrease in production rate and will maintain the reserves.